1) In a local bar, a customer slides an empty beer mug down the counter for a refill. The height of the counter is 1.16 m. The mug slides off the counter and strikes the floor 0.40 m from the base of the counter.

(a) With what velocity did the mug leave the counter? (.823) Correct: Your answer is correct. m/s

(b) What was the direction of the mug's velocity just before it hit the floor? Incorrect: Your answer is incorrect. How does the x component of the velocity of the mug just before it hits the floor compare to its initial velocity?° (below the horizontal)

2) A projectile is fired in such a way that its horizontal range is equal to 11.5 times its maximum height.

What is the angle of projection?
°
Answer #1

b)

time of flight

h = 0.5 g t^2

1.16 = 4.9* t^2

t = 0.487 s

vertical velocity of mug

vy = gt = 9.8* 0.487 = 4.768 m/s

Horizontal velocity, vx = 0.823 m/s

direction of mug velocity

phi = arctan ( vy/vx)

phi = 80.2

=========

2)

R = u^2 sin 2x / g

H = u^2 (sin x) ^2 / 2g

R = 11.5 H

u^2 sin 2x / g = 11.5 u^2 (sin x) ^2 / 2g

2 sin x cos x / g = 11.5 (sin x) ^2 / (2g)

4 cos x = 11.5 sin x

tan x = 4 / 11.5

x = 19.18

========

Comment in case any doubt, will reply for sure.. Goodluck